# PLAME: Lightweight MSA Design Advances Protein Folding From Evolutionary Embeddings

### v2

@@ 3.4 MSA Selection Method – HiFiAD
Existing MSA generation models struggle in ensuring the quality of generated MSAs. Different methods capture MSA distributions with varying effectiveness, and excessive concatenation of virtual MSAs can degrade the quality of the original MSA. This, in turn, restricts the ability of protein structure prediction software to interpret evolutionary distributions. While MSAGPT has explored MSA selection in its ablation studies, existing approaches lack a clear definition of MSA quality and fail to propose systematic selection methods based on sequence-level quality metrics chen2024msagpt .
During MSA augmentation, virtual MSAs inevitably introduce noise, limiting the performance of folding software. A common issue with current models is the generation of over-conserved sequences that closely resemble the query sequence. When these sequences are over-concatenated, they distort the evolutionary distribution of the original MSA. This issue was highlighted in MSAGPT’s studies, which showed that performance plateaued once the number of generated MSAs exceeded 32. To address this, balancing recovery rate distributions by selecting sequences with both high and low recovery rates has proven more effective than focusing exclusively on one or the other.
To address these issues, we propose HiFiAD to balance fidelity and diversity for MSA selection while maintaining efficiency. For orphan proteins without original MSAs (termed zero-shot cases), we select the top-$k$ sequences with the highest $S_{\text{BLOSUM}}$, along with sequences from the top and bottom $k/2$ of the recovery rate distribution. The BLOSUM substitution score is calculated as:
@@
where $B$ is the BLOSUM matrix (BLOSUM62 in our case); $s_{j}$ represents the j-th amino acid of sequence S with length L, and $m_{ij}$ refers to the j-th position of the i-th MSA sequence with length L. For proteins with existing MSAs (termed few-shot cases), we limit the total number of concatenated MSAs to reduce noise in the original MSA distribution. Specifically, we cap the number of MSA sequences $N_{\text{aug}_{\text{max}}}$ at $\max(16,2N)$. If fewer than $N_{\text{aug}_{\text{max}}}$ virtual MSAs are available, we retain all generated MSAs to ensure sufficient diversity for downstream analysis.
@@
pLDDT GDT TMscore RMSD LDDT pTM
@@
PLAME 71.50 70.48 0.58 0.59 0.58 0.64 11.41 10.62 0.60 0.64 / /
4 Experiment
Baselines
To evaluate PLAME’s capability in generating high-fidelity and diverse MSAs, we compared it with several state-of-the-art AI-based MSA generation methods and AlphaFold2’s MSA pipeline jumper2021alphafold . The baselines include AF2 MSA, and open-source methods including EvoDiff and MSAGPT chen2024msagpt ; alamdari2023protein . Additionally, to assess the potential advantages of leveraging AlphaFold2’s knowledge base in protein structure prediction, we included ESMFold lin2023evolutionary , an MSA-free folding software, as reference.
Evaluation Metric
We apply the following metrics for structure assessment: 1) local metrics including pLDDT and LDDT; 2) global metrics including GDT, TM-Score, pTM, and RMSD. To comprehensively evaluate the effectiveness of augmented MSAs, we tested three AF2 folding mode, progressing from computationally efficient, basic settings to more comprehensive configurations: 1) Mode1: Folding using pTM-3 model without using templates; 2) Mode2: Folding using the default 5 models without using templates; 3) Mode3: Folding using the default 5 models with templates. To evaluate the quality of generated multiple sequence alignments (MSAs), we employ four sequence-based metrics to quantify alignment fidelity and diversity: 1) Conservation Score measures the degree of residue conservation at each alignment position. It is defined as $C_{i}=\frac{\text{Freq}_{\text{max}}(i)}{N}$, where $\text{Freq}_{\text{max}}(i)$ is the frequency of the most common residue at position $i$, and $N$ is the total number of sequences. A high score indicates strong conservation, often reflecting functional or evolutionary constraints, while a low score suggests positional variability. 2) Gap Proportion quantifies the fraction of gaps at each alignment position, calculated as $G_{i}=\frac{G(i)}{N}$, where $G(i)$ is the number of gaps at position $i$. Lower values indicate better alignment quality with fewer missing data or alignment errors. 3) Substitution Compatibility assesses the evolutionary plausibility of aligned residues using BLOSUM62 substitution scores $S_{\text{BLOSUM}}$ (Eq. 16). Higher scores indicate evolutionary compatibility and alignment biological relevance. 4) Alignment Entropy captures residue diversity at each position using Shannon entropy:
$H_{i}=-\sum_{r\in\{R_{i}\}}p(r)\log_{2}p(r),$
where $\{R_{i}\}$ is the set of unique residues at position $i$, and $p(r)=\frac{\text{count}(r)}{N}$ is the relative frequency of residue $r$. Higher entropy signifies greater positional diversity, indicating weaker functional constraints, whereas lower entropy denotes stronger conservation, often associated with functional or evolutionary significance.
Datasets
For the training dataset, we use the PDB and UniClust30 subsets from the open-source OpenProteinSet as our data source ahdritz2024openproteinset . The pre-searched MSAs from OpenFold training are also included. We only retain data with at least 64 MSA sequences. To avoid overlap with the test cases, we removed sequences with over 90% similarity using MMSeqs based on UniClust30 clustering results mirdita2017uniclust ; steinegger2017mmseqs2 . This process yields an initial dataset of 293,979 samples, which are split into training and validation sets with a 90:10 ratio. For the test dataset, we adopt the curated test cases from MSAGPT chen2024msagpt , which consist of 200 protein samples from three benchmarks: CASP14&15, CAMEO haas2018continuous , and PDB berman2000protein . Any $90\%$ redundancy between the test cases and training dataset is eliminated.
4.1 PLAME as High-quality MSA Designer
General folding performance
We conducted experiments across three AF2 folding modes and evaluated six folding-related metrics. Results are shown in Table 3.4. We analyze the quality of MSAs generated by different methods and their performance in various scenarios. 1) Performance Superiority: PLAME achieved state-of-the-art performance on most metrics across all modes. This demonstrates that PLAME-generated MSAs effectively enhance AF2 folding, particularly in scenarios involving low-homology or orphan proteins, where significant improvements in prediction accuracy were observed. In contrast, EvoDiff and MSAGPT struggled to produce high-quality MSAs under these conditions. By leveraging the evolutionary latent space provided by ESM-2, PLAME generated biologically meaningful virtual MSAs, outperforming baselines on key metrics such as pLDDT and RMSD. 2) Diminishing Returns: As more advanced folding configurations were employed (e.g., with more powerful AF2 models or the inclusion of templates), the benefits of MSA augmentation gradually decreased. In Mode3, where templates were used, certain metrics exhibit performance decrease. This is likely because template information already captures much of the necessary sequence and evolutionary information, reducing the marginal utility of additional MSA augmentation. Furthermore, the noise introduced by low-quality MSAs can be amplified in stronger baselines, further impacting performance. 3) Metric Discrepancies: The enhancement effects across different metrics were not entirely consistent. For instance, the trends for pLDDT and RMSD diverged. While pLDDT, as a confidence metric, reflects the model’s confidence in its predictions, RMSD measures the global deviation of the predicted structure from the ground truth. MSA augmentation improved local prediction quality (leading to higher pLDDT), but had limited impact on reducing global structural deviation (resulting in minimal RMSD improvement). Nevertheless, pLDDT alone is sufficient to demonstrate the high quality of the generated MSAs. 4) Noise Issues: EvoDiff and MSAGPT produced MSAs that consistently underperformed compared to the original AF2 MSAs across all modes. This was particularly evident in low-homology protein scenarios, where the generated MSAs failed to accurately capture evolutionary distributions and often included irrelevant sequences. When concatenated with the original AF2 MSAs, these low-quality sequences introduced additional noise, disrupting the model’s inference. This aligns with the hypothesis presented in Section 3.4, where low-quality MSAs were predicted to introduce noise, ultimately impairing folding performance. 5) Few/Zero-shot Consistency: In Few-shot scenarios, EvoDiff and MSAGPT performed relatively better due to the presence of initial MSAs that provided additional homology information to guide the generation process. In contrast, PLAME does not rely on initial MSA searches and directly generates high-quality MSAs through its evolutionary latent space. This enables PLAME to excel in Zero-shot scenarios while maintaining consistent performance in Few-shot cases. This advantage significantly reduces runtime costs while ensuring robust enhancements across all scenarios.
Sequence quality assessment
@@ 4.1 PLAME as High-quality MSA Designer
To evaluate generated MSA quality, we designed a set of metrics focusing on MSA fidelity and diversity. These metrics aim to investigate which factors correlate most strongly with protein folding performance, especially given the lack of standardized criteria for such evaluations. Our study provides an initial exploration into this problem. Figure 3 illustrates the evaluation results.
PLAME-generated MSAs outperform other methods on fidelity metrics, with distributions closest to AF2 MSAs in Conservation Score, Gap Proportion, and Substitution Compatibility. Specifically, PLAME achieves higher Conservation Scores and Substitution Compatibility, reflecting its ability to better capture evolutionary and functional constraints. Moreover, its lower Gap Proportion indicates higher alignment completeness, which can be attributed to the latent evolutionary space providing richer homology information. These results highlight PLAME’s superior alignment with the evolutionary constraints of the target protein.
We analyzed Alignment Entropy as a measure of diversity. While greater diversity is generally expected to enhance homologous information, our results show that PLAME’s diversity levels are closer to those of AF2 MSAs, rather than exceeding methods like EvoDiff. This supports our assumption in 3.4 that excessive diversity can introduce noise and diminish the information enrichment effect of generated MSAs. Therefore, achieving a balance between fidelity and moderate diversity is crucial for MSA generation models that struggle to capture high-quality distributions.
4.2 PLAME as MSA-free ESMFold Adapter to AlphaFold2
The comparative analysis between PLAME and ESM-Fold demonstrates PLAME’s growing advantage as the folding modes become more advanced. In Mode1 and Mode2, PLAME outperforms ESMFold in key metrics such as pLDDT and RMSD, with obvious gains in pLDDT. In Mode3, the introduction of structural templates amplifies PLAME’s advantage, achieving better results across all metrics. For instance, pLDDT of PLAME increases from 66.54 to 71.50, surpassing ESMFold’s 66.26. These results validate PLAME’s ability to merge ESMFold’s computational efficiency with AF2’s performance by generating high-quality virtual MSAs. Acting as an adaptive network, PLAME transforms ESMFold’s single-sequence inference into AF2’s MSA-based folding, integrating evolutionary information to enhance prediction accuracy. Thus, PLAME overcomes ESMFold’s limitations by introducing virtual MSA generation, achieving significant improvements in Mode3 and providing an efficient and accurate solution for protein structure prediction.
Table 2: Ablation study over different MSA selection approaches.
pLDDT GDT TMscore RMSD LDDT pTM
@@ 4.2 PLAME as MSA-free ESMFold Adapter to AlphaFold2
HiFiAD 66.54 66.08 0.53 0.54 0.53 0.58 11.48 12.14 0.57 0.60 0.49 0.52
4.3 Ablation on Selection Methods
To evaluate the impact of MSA quality (3.4), we perform an ablation study using different MSA selection strategies. The results are summarized in Table 2. The selection methods are based on sequence similarity metrics (Rec) and BLOSUM substitution scores (BLOSUM). Additionally, Top and Down refer to the highest- and lowest-scoring cases. The experiments lead to the following key observations. 1) Effective HiFiAD filtering: HiFiAD achieves superior performance across all metrics compared to other filtering methods, demonstrating that the combination of high fidelity and balanced diversity is essential. 2) High-fidelity: Compared to the Top-down-Rec method and random selection, HiFiAD identifies more high-fidelity cases, validating the importance of leveraging in-distribution MSAs. 3) Diverse samples: Compared to BLOSUM-based methods and Top-Rec, HiFiAD selects more diverse samples. This balanced diversity prevents the conservation distribution from being overly deterministic, which could potentially correct the augmented MSA distribution. Among sequence-metric-based filtering methods, HiFiAD dynamically adjusts to Few- and Zero-shot scenarios, as well as varying MSA quality levels, making it a simple yet effective approach for MSA selection. More ablation results are shown in Table 9 in Appendix.
@@ 5 Conclusion
In this work, we introduced PLAME, the first model to leverage evolutionary embeddings for MSA sequence generation. Our approach bridges the gap between single-sequence inference and MSA-based methods, effectively enhancing protein folding performance. Our extensive evaluations demonstrate that PLAME-generated MSAs surpass existing methods in both conservation and diversity metrics, leading to significant improvements in structure prediction accuracy across various protein families. By functioning as both an MSA augmenter and an efficient AlphaFold adapter without the need for time-intensive MSA searches, PLAME establishes itself as a fast, accurate, and scalable solution for protein structure prediction. Furthermore, our proposed quality metrics and comprehensive experiments provide new insights into the relationship between MSA characteristics and folding performance, advancing our understanding of how sequence information translates to structural accuracy.

### v3

@@ 3.4 MSA Selection Method – HiFiAD
Generated MSAs often contain noise that degrades folding performance. We propose HiFiAD to balance fidelity and diversity in MSA selection through principled quality metrics.
HiFiAD addresses two key challenges: (i) over-conserved sequences that distort evolutionary distributions when over-concatenated, and (ii) lack of systematic quality assessment for generated MSAs, by combining sequence similarity (fidelity) with diversity to maintain balanced evolutionary signals.
Given a query sequence $s$ and generated MSAs $M=\{m_{1},m_{2},\ldots,m_{n}\}$, we define:
@@
$R(m_{i},s)$ $=\frac{1}{L}\sum_{j=1}^{L}\mathbb{I}[s_{j}=m_{ij}],\quad\forall m_{i}\in M,$ (16)
where $B$ is the BLOSUM62 matrix, $R(m_{i},s)$ is the recovery rate, and $\mathbb{I}[\cdot]$ is the indicator function.
Zero-shot selection (Orphan proteins): Select top-$k$ sequences by $S_{\text{BLOSUM}}$ and sequences from top/bottom $k/2$ of recovery rate distribution, similar to the Static Diversity Strategy of MSAGPT.
Few-shot selection (Low homology proteins): Limit augmented MSAs to $N_{\text{max}}=\max(16,2N_{\text{orig}})$ where $N_{\text{orig}}$ is the original MSA count. This design prevents evolutionary information distortion caused by excessive generated MSAs.
4 Experiment
Baselines
To evaluate PLAME’s capability in generating high-fidelity and diverse MSAs, we compared it with several state-of-the-art AI-based MSA generation methods and AlphaFold2’s MSA pipeline (jumper2021alphafold, ). The baselines include AF2 MSA (johnson2010hidden, ), and open-source methods including EvoDiff and MSAGPT (chen2024msagpt, ; alamdari2023protein, ). Additionally, we include an MSA-free method, ESMFold (lin2023evolutionary, ), to evaluate the complementary benefits of explicit MSA enhancement versus implicit evolutionary modeling.
Datasets
For the training dataset, we used the PDB and UniClust30 subsets from the OpenProteinSet as our data source (ahdritz2024openproteinset, ). The pre-searched MSAs from OpenFold training were also included. We retained data with at least 64 MSA sequences. To avoid overlap with the test cases, we removed sequences with over 90% similarity by MMSeqs based on UniClust30 clustering results (mirdita2017uniclust, ; steinegger2017mmseqs2, ). This process yielded an initial dataset of 293,979 samples, which were split into training and validation sets with a 90:10 ratio. For the test dataset, we adopted the curated test cases from MSAGPT (chen2024msagpt, ), which consist of 200 protein samples from three benchmarks: CASP14&15, CAMEO (haas2018continuous, ), and PDB (berman2000protein, ). Any $90\%$ redundancy between the test cases and training dataset was eliminated.
Evaluation
Structural Assessment Metric We evaluate structure qualitywith local and global metrics. Local metrics include pLDDT (per-residue confidence) and LDDT (local distance difference test). Global metrics comprise GDT (global distance test), TM-Score (template modeling score) (zhang2005tm, ), pTM (predicted TM-score), and RMSD (root mean square deviation).
AlphaFold2 Folding Modes To comprehensively assess MSA augmentation effectiveness, we evaluate three AF2 configurations with increasing computational complexity:
Sequence Assessment Metric We employ four sequence-based metrics to quantify alignment fidelity and diversity:
1) Conservation Score measures residue conservation at each position: $C_{i}=\text{Freq}_{\text{max}}(i)/N$, where $\text{Freq}_{\text{max}}(i)$ is the most frequent residue at position $i$ and $N$ is the sequence count. Higher scores indicate stronger evolutionary constraints.
2) Gap Proportion quantifies alignment completeness: $G_{i}=G(i)/N$, where $G(i)$ counts gaps at position $i$. Lower values indicate better alignment quality.
3) Substitution Compatibility evaluates evolutionary plausibility using BLOSUM62 scores $S_{\text{BLOSUM}}$ (Eq. 15). Higher scores reflect greater biological relevance.
4) Alignment Entropy captures positional diversity via Shannon entropy:
$H_{i}=-\sum_{r\in\{R_{i}\}}p(r)\log_{2}p(r)$ (17)
where $\{R_{i}\}$ represents unique residues at position $i$ and $p(r)=\text{count}(r)/N$. Higher entropy indicates greater diversity; lower entropy suggests functional conservation.
@@ 4 Experiment
pLDDT  ($\uparrow$) GDT  ($\uparrow$) TMscore  ($\uparrow$) RMSD($\downarrow$) LDDT  ($\uparrow$) pTM  ($\uparrow$)
@@
[…]
AF2 MSA 66.34 72.54 0.55 0.61 0.56 0.65 11.29 10.29 0.58 0.66 / /
[…]
4.1 Structure Benchmark Assessment
We evaluated PLAME across three AF2 folding modes and AF3, using six structural metrics to assess MSA generation quality (See details in Table4).
4.1.1 General Performance Comparison
PLAME demonstrates consistent superiority across both zero-shot and few-shot scenarios against traditional MSA searching, AI-based searching, and AI-based generative methods, establishing a new paradigm for MSA generation without traditional homology search. In zero-shot settings, where proteins lack existing MSAs, PLAME achieves remarkable improvements with pLDDT scores reaching 71.50 in Mode3, significantly outperforming competing methods like EvoDiff (64.39) and MSAGPT (68.39). Moreover, the performance gap becomes even more pronounced in challenging scenarios: while EvoDiff and MSAGPT often introduce detrimental noise when their generated sequences are concatenated with original AF2 MSAs, PLAME consistently enhances folding quality. Interestingly, few-shot scenarios reveal that existing methods can partially recover performance when guided by initial homologous sequences, yet PLAME maintains its edge by generating more coherent evolutionary profiles that complement rather than interfere with existing MSAs.
4.1.2 Mode-Dependent Performance Patterns
The progression from Mode1 through AF3 reveals intriguing insights about the relationship between model sophistication and MSA augmentation benefits. Mode1 and Mode2 demonstrate the strongest relative improvements from PLAME-generated MSAs, with pLDDT gains of up to 5 points across different baseline methods. As configurations advance to Mode3 with structural templates, the enhancement effects become more nuanced—while absolute performance continues to improve, the marginal gains from MSA augmentation diminish because template information already captures substantial evolutionary constraints. This phenomenon reflects a fundamental trade-off in modern protein folding: as models become more powerful and incorporate diverse information sources, the additional value of synthetic MSAs decreases, though PLAME’s high-quality generations continue to provide meaningful contributions. The AF3 results further validate this trend, showing that PLAME maintains its effectiveness even with more advanced folding architectures, suggesting that high-quality virtual MSAs remain valuable complements to cutting-edge structural prediction methods.
4.1.3 PLAME vs ESMFold: Bridging Efficiency and Accuracy
The comparison with ESMFold reveals PLAME’s unique position in the protein folding landscape, offering a compelling alternative that combines computational efficiency with enhanced accuracy. While ESMFold achieves reasonable baseline performance (pLDDT of 66.26), PLAME progressively widens this gap as more sophisticated folding configurations are employed. In basic Mode1, PLAME shows modest improvements, but the advantage becomes substantial in Mode3 where PLAME reaches 71.50 pLDDT compared to ESMFold’s unchanged 66.26. This trend suggests that PLAME-generated MSAs provide increasingly valuable evolutionary context that more advanced folding models can effectively exploit. The consistent RMSD improvements across all modes further validate that PLAME’s virtual MSAs contribute meaningful structural constraints, enabling users to achieve AF2-level accuracy while maintaining the computational advantages of MSA-free approaches.
4.2 Sequence quality assessment
To evaluate generated MSA quality beyond structural perspectives, we conducted sequence-level analysis by fidelity and diversity metrics. This provides an additional critical gap—establishing criteria for understanding generated MSA quality. Figure 3 presents our comparative analysis.
PLAME achieves superior evolutionary fidelity by closely mimicking the distributional characteristics of natural MSAs across all key metrics. The results reveal PLAME’s distributions align most closely with AF2 MSAs in Conservation Score, Gap Proportion, and Substitution Compatibility, demonstrating its ability to capture authentic evolutionary constraints.
@@ 4.2 Sequence quality assessment
This fidelity advantage manifests in higher Conservation Scores and Substitution Compatibility values, indicating that PLAME-generated sequences preserve functionally critical residues while incorporating biologically plausible substitutions. The significantly lower Gap Proportion validates PLAME’s approach, as the evolutionary latent space from ESM-2 provides richer homology information enabling more complete alignments.
PLAME maintains diversity levels comparable to natural AF2 MSAs, supporting our hypothesis that excessive diversity introduces detrimental noise. Rather than maximizing diversity like EvoDiff, this measured approach aligns with our selection strategy principles, where balanced information enrichment proves more effective than naive sequence proliferation (Section 3.4). The findings suggest successful MSA generation requires maintaining the delicate balance between providing sufficient homologous information and avoiding noise from unconstrained sequence generation, positioning PLAME as a method that respects fundamental biological constraints.
Table 2: Ablation study of HiFiAD on PLAME and other baselines.
pLDDT  ($\uparrow$) GDT($\uparrow$) TMscore  ($\uparrow$) RMSD  ($\downarrow$) LDDT  ($\uparrow$) pTM  ($\uparrow$)
@@
EvoDiff-HiFiAD 58.24 60.89 0.46 0.49 0.46 0.54 13.74 12.39 0.51 0.56 / /
MSAGPT-HiFiAD 60.16 62.63 0.48 0.52 0.48 0.57 12.54 12.18 0.53 0.59 / /
DHR-HiFiAD 66.01 66.08 0.53 0.55 0.53 0.60 11.48 12.14 0.57 0.60 / /
[…]
4.3 Ablation Studies
To validate our HiFiAD selection strategy, we conducted ablation experiments across different selection approaches and baseline methods. Table 2 compares various selection strategies and evaluates HiFiAD’s effectiveness on other generative methods.
HiFiAD consistently outperforms alternative selection strategies by optimally balancing fidelity and diversity constraints. Compared to similarity-based methods (Top/Down-Rec) and substitution matrix approaches (BLOSUM-32), HiFiAD achieves superior performance with pLDDT scores of 66.54 in zero-shot settings, demonstrating the importance of jointly considering evolutionary fidelity and controlled diversity. The strategy effectively identifies high-fidelity sequences while maintaining sufficient diversity to prevent overly deterministic conservation patterns. HiFiAD automatically adapts to varying MSA quality levels and shot configurations, making it robust without requiring manual parameter tuning.
When applied to competing baselines, HiFiAD consistently improves performance: EvoDiff benefits from a 58.24 to 60.89 pLDDT improvement, MSAGPT gains from 60.16 to 62.63, and DHR advances from 66.01 to 66.08. These improvements demonstrate that HiFiAD addresses fundamental challenges in MSA selection across all generative approaches. The consistent gains across different generation paradigms—from diffusion-based (EvoDiff) to autoregressive (MSAGPT) and retrieval-based (DHR) methods—validate that the fidelity-diversity trade-off represents a universal principle in MSA augmentation. The improvement margins correlate with baseline method quality, suggesting HiFiAD provides proportional benefits while maintaining relative performance hierarchy.
Also, we conducted ablation study on MSA length (See Table3). PLAME shows overall improvement on all length ranges, where performs the largest improvement on 100-300 range. We believe this is because the MSA training data are mainly concentrated in this range (chen2024msagpt, ).
Table 3: Ablation on protein length.
Length Range pLDDT($\uparrow$) GDT($\uparrow$) TMscore($\uparrow$) RMSD($\downarrow$) LDDT($\uparrow$)
AF2 MSA <<100 71.03 0.64 0.52 7.77 0.61
AF2 MSA 100-300 59.50 0.49 0.53 12.46 0.54
AF2 MSA >>300 56.29 0.43 0.51 15.67 0.53
PLAME <<100 74.12 0.63 0.52 7.49 0.61
PLAME 100-300 65.55 0.53 0.58 11.58 0.58
PLAME >>300 58.31 0.45 0.53 16.16 0.54
Furthermore, we provide additional case studies on folding enhancement. More case studies on orphan de novo proteins (SectionE.5), protein failure cases (SectionE.4), selected protein cases with aligned structures (SectionE) can be found in the appendix.
Figure 4: Case study of folding enhancement of PLAME, MSAGPT, and AF2 MSA on 8okh_B.
@@ 5 Conclusion
In this study, we introduce PLAME, the first model to leverage evolutionary embeddings for MSA generation and apply it to protein folding enhancement. Our approach bridges the gap between single-sequence inference and MSA-based methods, effectively improving protein folding performance. Evaluation results demonstrate that PLAME-generated MSAs outperform existing methods in both conservation and diversity metrics, significantly enhancing structural prediction accuracy across different protein families. PLAME serves as both an MSA enhancer and an efficient AlphaFold adapter without requiring time-consuming MSA searches, providing a fast, accurate, and scalable protein structure prediction solution. Additionally, our proposed quality metrics and experiments offer new insights into the relationship between MSA features and folding performance.